OpenStudy (anonymous):

Please help! Add, subtract, multiply, or divide as indicated. Make sure that the final answer is simplified completely. 2/x-2*x^2-4/4

OpenStudy (anonymous):

order of addition- PEMDAS

OpenStudy (dumbcow):

\[\frac{2}{x-2}*\frac{x^{2}-4}{4}\]

OpenStudy (dumbcow):

factor x^2 -4

OpenStudy (anonymous):

diff of 2 perfect squares

OpenStudy (dumbcow):

\[x^{2}-4 = ?\]

OpenStudy (anonymous):

-4

OpenStudy (dumbcow):

don't think number answer, you are factoring into (x+a)(x+b) form

OpenStudy (dumbcow):

a*b = -4 a+b = 0 what is a, what is b

OpenStudy (dumbcow):

any ideas? what are 2 numbers that multiply to -4

OpenStudy (anonymous):

-4, 1?

OpenStudy (dumbcow):

yep now check to see if they add up to 0

OpenStudy (anonymous):

-3, so no.

OpenStudy (dumbcow):

so try 2 other numbers that multiply to -4

OpenStudy (anonymous):

-2, 2

OpenStudy (anonymous):

x-2/2?

OpenStudy (dumbcow):

sorry i got cut off from my computer... so a and b are -2,2 \[x^{2}-4 = (x+2)(x-2)\] Now you can cancel and reduce

OpenStudy (dumbcow):

oh i see..almost the (x-2) is on the bottom so that is what cancels leaving (x+2) on top answer: (x+2)/2